A rod in a steam jacket is measured to have a length of 0.600m at a temperature of 20.0 degrees C. Steam is then passed through the jacket for several minutes until the rod is at a temperature of 98.0 degrees C. The increase in the length of the rod is measured by a micrometer screw arrangement to be 1.19mm. What is the linear thermal coefficient of expansion a for the rod?
Answer #1

initial length of the rod l = 0.6 m

initial temperatuire T1 = 20 degreess

final length l _o = 0.6 + 0.00119 = 0.60119 m

final temperature of the rod   T2 = 98 degrees

       L = Lo ( 1 - α ΔT)

        0.60119 = 0.6 ( 1 + α 78 )

     the linear thermal coefficient of thermal expansion is    

              α = 2.54*10^ -5 C o ^ ( -1 )